OpenStudy (drummergirl7):

In 2/5 hour, Matt can type 4/5 page. What is Matt's rate in pages per hour? Drag and drop the correct value into the box. 8/25 3 1/8 1/2 2

OpenStudy (drummergirl7):

@steve816

OpenStudy (aveline):

Multiply each side of the equation by 5/2

OpenStudy (drummergirl7):

hi :)

OpenStudy (drummergirl7):

how do i do that? sorry im really bad at math :( do i multiply both numbers by 5/2

OpenStudy (shadowlegendx):

\[\frac{ 2 }{ 5 }h = \frac{ 4 }{ 5 }p\] Where h = an hour and p = a page

OpenStudy (drummergirl7):

i think its c but i could be wrong what do u think??

OpenStudy (shadowlegendx):

Multiply both sides by 5/2 \[\frac{ 5 }{ 2 }(\frac{ 2 }{ 5 }h) = \frac{ 5 }{ 2 } (\frac{ 4 }{ 5}p)\] \[\frac{ 10}{ 10 }h = \frac{ 20 }{ 10}p\] 10/10 = 1 20/10 = 2 \[1h = 2p \rightarrow h = 2p\] 1 hour = 2 pages

OpenStudy (drummergirl7):

so its 1/2 :D

OpenStudy (shadowlegendx):

How

OpenStudy (drummergirl7):

u said 1 hour 2 pages does that equal 1/2?

OpenStudy (shadowlegendx):

If it's saying \[\frac{ 1hour }{ 2pages }\] Then yeah

OpenStudy (drummergirl7):

yay thx so much

OpenStudy (shadowlegendx):

No problem